OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL! Luis starts at one corner of a square field and walks 80 feet along one side to another corner of the field. He turns 90° and walks 60 feet, and then walks straight back to where he started. What is the area of the part of the field he walked around? A. 140 ft² B. 280 ft² C. 320 ft² D. 2400 ft²

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

Dont give direct answers please. I know its unavoidable sometimes, but it could get us both banned

OpenStudy (anonymous):

 Use pythogoras theorem, \[c^2 = a^2 + b^2\]

OpenStudy (anonymous):

and I meant D

OpenStudy (anonymous):

sorry

OpenStudy (faiqraees):

yes it should be d

OpenStudy (anonymous):

Again, @baljeet16, youre speaking a whole different language to me.

OpenStudy (anonymous):

Ohh sorry you need to find area so use Area = \[\frac{1}{2}ab\]

OpenStudy (faiqraees):

yes 1/2 * 80*60

OpenStudy (anonymous):

Can you help me with this one...

OpenStudy (anonymous):

multiply the two numbers and divide by 2

OpenStudy (anonymous):

A farmer was counting his cows and chickens. He counted 13 heads and 36 legs. How many cows does he have? A. 5 B. 8 C. 9 D. 18

OpenStudy (anonymous):

36 divided by 4 =

OpenStudy (faiqraees):

x+y =13 2x+4y=36 x is chicken y is cow

OpenStudy (anonymous):

cows has 4 legs and chicken 2. x + y = 13, where x is no of cow, and y= no of chicken 4*x + 2.y = 36

OpenStudy (faiqraees):

solve the simultaneous equation

OpenStudy (anonymous):

36 divided by 4 is 9

OpenStudy (anonymous):

oh hes counting chickens to :P

OpenStudy (anonymous):

Yeah, Im just gonna go with 9 as my answer

OpenStudy (anonymous):

And my grade is...

OpenStudy (anonymous):

80%, barely passed it!!!

OpenStudy (anonymous):

ha GG

OpenStudy (anonymous):

If you say so lol

OpenStudy (anonymous):

9 was wrong anwer.

OpenStudy (anonymous):

I dont care, at least I passed

OpenStudy (anonymous):

answer is 5 cows and 8 chickens.

OpenStudy (faiqraees):

5 was correct

OpenStudy (anonymous):

Ok bye